OpenStudy (anonymous):

A certain function f has the property that f(x+y)= f(x)+f(y) for all values of x and y. Which of the following statements must be true when a=b? I. f(a+b)=2f(a). II. f(a+b)=[f(a] to the power of 2. III. f(b) +f(b)=f(2a)

OpenStudy (anonymous):

this is a function problem...

OpenStudy (cruffo):

(i) is true f(a+b) = f(a) + f(b) = f(a) + f(a) = 2f(a)

OpenStudy (anonymous):

f(a+b) = f(a+a) = f(a) + f(a) = 2f(a)

OpenStudy (anonymous):

ahh, since a=b, right?

OpenStudy (anonymous):

I. f(a+b)=2f(a) is true because f(a+b) = f(a) + f(b), since a = b, f(b) = f(a), so f(a) + f(a) = 2f(a) so I must be true.

OpenStudy (anonymous):

all good responses

OpenStudy (cruffo):

(iii) is also true. Just use what is given.

OpenStudy (anonymous):

let us see

OpenStudy (anonymous):

II. isn't true

OpenStudy (anonymous):

And because of this, II is false.

OpenStudy (anonymous):

a=1, b=1

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

a=3

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

6 does not equal 9

OpenStudy (anonymous):

but f(2a)

OpenStudy (anonymous):

because

OpenStudy (anonymous):

f(b)+f(b)=f(2b)=f(2a)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

ty micah

OpenStudy (cruffo):

f(b)+f(b)=f(2b)=f(2a) is a good statement. I'm not sure why you are using #'s like 1 and 3

OpenStudy (anonymous):

sometimes I like to pick variables to solve the problem...

OpenStudy (cruffo):

Usually, that's a good strategy, but you don't know what f(x) is in this case.

OpenStudy (anonymous):

We don't know function so you can't use number as a variable. We only know that f(a + b) = f(a)+ f(b), that all.

OpenStudy (anonymous):

Just like cruffo said.

OpenStudy (anonymous):

ty, guys

OpenStudy (anonymous):

next problem?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ask as many questions as you like.

OpenStudy (anonymous):

ty